The area of a square with side s is A(s) = s
2
.
What is the rate of change of the area of a square with respect
to its side length when s = 3
1
Expert's answer
2020-03-10T10:10:28-0400

"A(s)=s^2"

The rate change of area with respect to side is given by

"\frac{d}{ds}(A(s))=2s"

And "s=3"

"\frac{d}{ds}(A(s))=2s=2 \times 3=6m^2/m"